Number of Inquiries (X) = 22 , 23 , 24 , 25 , 26 , 27 Probability P(X) = 0.08, 0.19, 0.36, 0.25, 0.07, 0.05 What is the mean? *

1
Expert's answer
2022-04-25T16:09:50-0400

"\mu=\sum X\cdot P(X)=22\cdot 0.08+23\cdot 0.19+24\cdot 0.36+25\cdot 0.25+26\cdot 0.07+27\cdot 0.05=24.19"


Answer: "\mu=24.19"